# A Hybrid Quantum Neural Network for Split Learning

### v1

@@ 6.3.2 Inference Time Performance of HQSL and SL At Different Noise Levels
From Section 6.3.1, we established that when $\mu=2\pi$ and $4\pi$, we obtain large differences between original and reconstructed images indicated by the deviations above the baseline results. In this section, we investigate the effect of varying scale parameters, $b$, on the inference performance of HQSL and SL given mean values $\mu=2\pi$ and $4\pi$. We measure the performance by comparing the accuracy and F1-score in classifying FMNIST, MNIST, and Speech Commands spectrogram images with varying scale parameters, $b$. We present these results using the box plots shown in Fig. 13 and 14, representing the five-fold test accuracies and F1-scores in the presence and absence of noise. The baseline performances are represented by the Noise-free box plots.
@@
Hence, we conclude here that, by tuning the Laplacian noise layer parameters considering the rotational properties of encoding gates, HQSL’s classification performance is as high as its baseline performance, showing robustness to the noise layer. In contrast, SL is less robust to the tuned noise layer.
@@
Fig. 15 illustrates that the reconstructed images show more significant differences from the original images in the hybrid setting than in the classical setting. Hence, we can directly confirm that our designed noise layer effectively impedes the reconstruction model in recreating the raw images. This is also depicted in the results shown in Fig. 12. Furthermore, in HQSL, the noise layer causes the smashed data to be handled properly by the quantum layer on the server-side model, hence maintaining a similar classification performance to the noise-free performance as shown in Fig. 13 and 14. This is due to the configured mean values of $\mu=2\pi$ and $4\pi$, which causes the quantum state encoded by the RX-gates in the quantum layer to be nearly invariant to the additive Laplacian noise, given the scale parameter is small ($b=0.01)$. This means that, for the quantum layer, the obfuscated classical intermediate features correspond to quantum states that are close to those encoded from clean classical smashed data.
By leveraging the rotational properties of the encoding gates in the quantum layer of HQSL, we can tune the noise layer parameters to obtain clear advantages in terms of enhancing security in split learning against reconstruction attacks from data privacy leakage. The advantages of our proposed noise layer in HQSL are that it (i) impedes the reconstruction attack model’s performance in reconstructing raw data, and (ii) enables HQSL to maintain a high performance despite its presence. HQSL addresses the privacy-utility tradeoff associated with the use of noise layers to enhance security against data privacy leakage in SL.
@@ 7 Conclusion
This paper addresses the problem of applying Split Learning (SL) concepts to pure Quantum Machine Learning (QML) models for resource-constrained clients lacking quantum computing resources. We propose a novel Hybrid Quantum Split Learning, HQSL, architecture that enables resource-limited clients in the classical domain to train their models collaboratively with a hybrid quantum server for classification tasks. Specifically, HQSL consists of a Hybrid Quantum Neural Network, HQNN, split into two parts: the client side, consisting of a classical neural network, and the server side, comprising an HQNN with the first layer as a quantum layer consisting of a 2-qubit quantum circuit. For the quantum circuit, we introduce a qubit-efficient data-loading technique that allows us to keep the number of qubits and circuit depth small in anticipation of their implementation on near-term quantum devices. Experimental results demonstrated the feasibility and ability of HQSL to improve classification accuracy and F1-score compared to its classical equivalent. Our experiments also demonstrated the scalability of HQSL with an increasing number of clients as HQSL maintains high performance during classification as we sequentially increase the number of clients.
@@
In this work, we also address the overarching problem of data privacy leakage associated with SL, which can lead to clients’ private input data reconstruction attacks on the server side. We propose a Laplacian noise layer defense mechanism designed based on the rotational properties of the encoding gates present within the quantum layer of HQSL. By tuning the parameters of our proposed Laplacian noise layer defense mechanism, we showed that HQSL is more robust to the privacy-utility tradeoff generally associated with the use of noise to enhance security against reconstruction attacks in SL.
@@
Overall, our novel HQSL enables a resource-constrained client to train its ML model collaboratively with a server equipped with quantum computing resources, resulting in improved classification performance. The quantum circuit we use in this work underscores a qubit-efficient data-loading mechanism that can be extended beyond the scope of HQSL to address qubit count limitations in near-term devices. Furthermore, HQSL supports multiple clients in the classical domain for collaborative HQNN model training. The defense mechanism proposed in this work enhances HQSL’s security against data privacy leakage and the risks of reconstruction attacks.
Although a small quantum circuit was employed in the design of HQSL, it lays the foundation for future research involving Hybrid QML and SL. Furthermore, as fault tolerance in quantum computers becomes feasible, HQSL could be scaled with more complex quantum circuits and potentially involve more than one quantum layer in the server-side model. This would allow resource-constrained clients to further benefit from the potential advantages of quantum computing. Future research could also explore implementations of HQSL models in configurations other than the vanilla split, such as the U-shaped configuration for SL without label sharing or SL using vertically partitioned data.

### v2

@@ 6.3.2 Inference Time Performance of HQSL and SL At Different Noise Levels
From Section 6.3.1, we established that when $\mu=2\pi$ and $4\pi$, we obtain large differences between original and reconstructed images indicated by the deviations above the baseline results. In this section, we investigate the effect of varying scale parameters, $b$, on the inference performance of HQSL and SL given mean values $\mu=2\pi$ and $4\pi$. We measure the performance by comparing the accuracy and F1-score in classifying FMNIST, MNIST, and Speech Commands spectrogram images with varying scale parameters, $b$. We present these results using the box plots shown in Fig. 14 and 15, representing the five-fold test accuracies and F1-scores in the presence and absence of noise. The baseline performances are represented by the Noise-free box plots.
@@
Hence, we conclude here that, by tuning the Laplacian noise layer parameters considering the rotational properties of encoding gates, HQSL’s classification performance can match the baseline performance, showing robustness to the noise layer. In contrast, SL is less robust to the tuned noise layer.
@@
Fig. 16 illustrates that the reconstructed images show more significant differences from the original images in the hybrid setting than in the classical setting. Hence, we can directly confirm that our designed noise layer effectively impedes the reconstruction model in recreating the raw images. This is also depicted in the results shown in Fig. 13. Furthermore, in HQSL, the noise layer causes the smashed data to be handled properly by the quantum layer on the server-side model, hence maintaining a similar classification performance to the noise-free performance as shown in Fig. 14 and 15. This is due to the configured mean values of $\mu=2\pi$ and $4\pi$, which causes the quantum state encoded by the RX-gates in the quantum layer to be nearly invariant to the additive Laplacian noise, given the scale parameter is small ($b=0.01)$. This means that, for the quantum layer, the obfuscated classical intermediate features correspond to quantum states that are close to those encoded from clean classical smashed data.
By leveraging the rotational properties of the encoding gates in the quantum layer of HQSL, we can tune the noise layer parameters to obtain clear advantages in terms of enhancing resistance in split learning against reconstruction attacks from data privacy leakage. The advantages of our proposed noise layer in HQSL are that it (i) impedes the reconstruction attack model’s performance in reconstructing raw data, and (ii) enables HQSL to maintain a high performance despite its presence. HQSL addresses the privacy-utility tradeoff associated with the use of noise layers to enhance resistance against data privacy leakage in SL.
@@ 7 Conclusion
This paper addresses the problem of applying Split Learning (SL) concepts to pure Quantum Machine Learning (QML) models for resource-constrained clients lacking quantum computing resources. We propose a novel Hybrid Quantum Split Learning, HQSL, architecture that enables resource-limited clients in the classical domain to train their models collaboratively with a hybrid quantum server for classification tasks. Specifically, HQSL consists of a Hybrid Quantum Neural Network, HQNN, split into two parts: the client side, consisting of a classical neural network, and the server side, comprising an HQNN with the first layer as a quantum layer consisting of a 2-qubit quantum circuit. For the quantum circuit, we introduce a qubit-efficient data-loading technique that allows us to keep the number of qubits and circuit depth small in anticipation of their implementation on near-term quantum devices. Experimental results demonstrated the feasibility and ability of HQSL to improve classification accuracy and F1-score compared to its classical equivalent. Performance evaluations on actual quantum devices and noisy simulators provide initial evidence supporting HQSL’s noise resilience and practicality for deployment on near-term, noisy quantum (NISQ) devices. Our experiments also demonstrated the scalability of HQSL with an increasing number of clients as HQSL maintains high performance during classification as we sequentially increase the number of clients.
@@
In this work, we also address the overarching problem of data privacy leakage associated with SL, which can lead to clients’ private input data reconstruction attacks on the server side. We propose a Laplacian noise layer defense mechanism designed based on the rotational properties of the encoding gates present within the quantum layer of HQSL. By tuning the parameters of our proposed Laplacian noise layer defense mechanism, we showed that HQSL is more robust to the privacy-utility tradeoff generally associated with the use of noise to enhance SL’s resistance against reconstruction attacks in SL.
@@
Overall, HQSL enables resource-constrained clients to train ML models collaboratively with a server equipped with quantum computing resources, resulting in improved classification performance. The quantum circuit designed in this work underscores a qubit-efficient data-loading mechanism that can be extended beyond the scope of HQSL to address qubit count limitations in near-term devices. Furthermore, HQSL supports multiple clients in the classical domain for collaborative HQNN model training. The defense mechanism proposed in this work enhances HQSL’s resistance against data privacy leakage and the risks of reconstruction attacks.
Future work. Although our current HQSL prototype uses a relatively small and shallow quantum circuit, it lays the groundwork for more ambitious and feasible HQNN+SL architectures. As fault‐tolerant devices become available, one could scale HQSL with richer variational blocks or multiple quantum layers on the server side, enabling even lightly resourced clients to benefit from potential quantum advantages. Future studies could also explore alternative split topologies, e.g., U-shaped splits without label sharing or vertically partitioned data, and benchmark HQSL against federated learning (FL) and quantum federated learning (QFL) baselines to provide a more comprehensive performance context. Moreover, while our experiments demonstrate that the rotationally-tuned Laplacian noise layer effectively preserves classification accuracy and impedes reconstruction attacks, a rigorous theoretical analysis, such as formal differential-privacy guarantees or robustness against adaptive adversaries, remains an open area for future research. In addition, future work could investigate defense techniques against a broader range of attack scenarios, beyond reconstruction attacks, e.g., model inversion, membership inference, and quantum-specific attacks, to deepen insights into HQSL’s security vulnerabilities and further strengthen its robustness.
@@ Appendix E Additional Experiments with Reconstruction Attack Models 2 and 3
Appendix F Speech Commands Dataset Result for Single Client-Single Server Experiments
In this section, we present the raw data collected for the Speech Commands dataset describing the accuracies and F1-scores obtained during the single client-single server experiments. The results presented in Table 8 complement the single client-single server results in Fig. 7 presented in Section 5.1. The data show that the extremely low standard deviations (std) in the accuracies and F1-scores for HQSL demonstrate that our HQSL model’s performance shows high consistency, corresponding to high stability on this dataset. These results also explain the narrow box plots obtained for the Speech Commands dataset using our proposed HQSL model, displayed in Fig. 7.
Table 8: Raw data for Speech Commands dataset for single client-single server experiments complementing Fig. 7. The low standard deviations (std) relative to the means for HQSL case demonstrate high consistency in the accuracy and F1-score results
Fold F1-Score Accuracy Loss
Classical (SL) 1 0.941733 0.941766 0.120393
2 0.951114 0.951158 0.105335
3 0.942385 0.942392 0.114520
4 0.951102 0.951158 0.105450
5 0.956093 0.956168 0.101465
mean 0.948485 0.948528 -
std 0.006214 0.006237 -
Hybrid (HQSL) 1 0.957421 0.957420 0.086337
2 0.959869 0.959925 0.082089
3 0.961175 0.961177 0.082415
4 0.959919 0.959925 0.086670
5 0.959918 0.959925 0.084178
mean 0.959660 0.959674 -
std 0.001368 0.001372 -